OpenStudy (anonymous):

Find the slope and y-intercept of the line. 18x + 4y = 112

OpenStudy (anonymous):

convert the equation in slope intercept form .

OpenStudy (anonymous):

y=mx + c where m is slope and c is the y intercept

OpenStudy (anonymous):

that doesn't help at all, dude.

OpenStudy (anonymous):

4y=-18x+112

OpenStudy (igreen):

@DefameMe Yes, it does. y = mx + b 'm' is the slope.

OpenStudy (anonymous):

y=(-9/2)x+112/4

OpenStudy (anonymous):

here slope is m =-9/2 and y intercept is 112/4

OpenStudy (anonymous):

hope u got it .